During the constant pressure reversible non-flow process with air, the temperature increases from 400 0F to 1600 0F. Show process on the p-V and T-s coordinates and find the heat transfer in BTU/lbm.
1
Expert's answer
2020-05-14T14:37:40-0400

Here, mass= 1 kg, Pressure = constant , "T_1=400 ^oF, T_2= 1600^0F"


Heat = "nC_p dT = 1\times 0.24\times (1600 - 400)=288" kJ